OpenStudy (anonymous):

The sum of two numbers is 36. One number is 2 times as large as the other. What are the numbers?

OpenStudy (jskhupmang):

the equation is this 2x+x=36

OpenStudy (jskhupmang):

do you know how to solve? @KiaraIsABro

OpenStudy (jskhupmang):

Well I don't time to wait for you so here how to calculate combine like terms 2x+x= 3x 3x= 36 divide both sides by 3 x=12 so since x equals 12 you can plug back into the equation 2x+x=36 24+12=36